# A Complex Scaling Method for Efficient and Accurate Scattering Emulation in Nuclear Reactions

### v1

@@ 1 Introduction
Nuclear reactions play a crucial role in helping us understand how nuclei are formed and the properties of dense nuclear matter, like that found in neutron stars [1]. These reactions are fundamental to many astrophysical events, especially those happening in stars and during explosive occurrences like supernovae. This makes them very important for the field of astrophysics.
However, the models we use to understand these reactions depend heavily on certain assumptions, like the optical model potential, which is a way to describe how particles interact. These models need parameters that are adjusted to fit the observed data, such as elastic scattering cross sections. This fitting process introduces a lot of uncertainties into our analysis of nuclear reactions. To deal with these uncertainties, it is vital to use robust methods for error quantification, such as Bayesian analysis, to make our results more reliable and understandable [2, 3].
@@
Several types of scattering emulators have been developed. For example, emulators based on the Kohn variational principle [6, 7], other variational methods [8, 9], and the $R$-matrix theory [10, 11] have shown impressive accuracy and efficiency. These emulators can quickly provide results that are very close to those obtained from detailed and time-consuming calculations.
However, there are some limitations. One major issue is that most emulators are designed to work with specific angular momentum states, which means they cannot easily be used for different angular momentum settings. This specialization increases the amount of storage needed and the computational time required, as the unique properties of each angular momentum state are not fully utilized. This limitation restricts the overall utility of these emulators, highlighting the need for more flexible and general approaches.
@@
The paper is organized as follows: In Sec. 2, we provide a brief overview on the complex scaling method and the development of our scattering emulator. Sec. 3 presents some application of our emulator, and an anomaly detection test. Finally, we conclude with a summary and discusses future research directions in Sec. 4.
@@ 2 Formalism
$U_{N}(r;\bm{\omega})=$ $V_{C}(r;R_{c})+V_{0}f\left(r,R_{R},a_{R}\right)$ (2)
$+iW_{0}f\left(r,R_{W},a_{W}\right)+iW_{S}\frac{\mathrm{d}}{\mathrm{d}r}f\left(r,R_{ws},a_{ws}\right),$
where $f\equiv(1+\mathrm{exp}((r-R)/a))^{-1}$ is the Woods-Saxon (WS) function. The parameter configuration is:
@@
where $\bm{\Omega}=\{\bm{\omega},\ell\}$ is the total parameter configuration we consider in this work. It should be noted that in some conventional methods for solving Eq. (1), the angular momentum quantum number $\ell$ serves not only as a parameter in the Hamiltonian but also specifies the boundary condition of the scattering wave function. This presents significant challenges when dealing with states of different angular momentum. To remove the dependence on different boundary conditions and treat $\ell$ just as a parameter, we apply the complex scaling method (CS) to solve Eq. (1). After separating the wave function and rotating the coordinate [12], the radial Schrödinger equation can be transformed into:
@@
where $\theta$ is the rotation angle in the complex scaling operation, $H^{\theta}$ is the complex-scaled Hamiltonian, $\psi_{\ell}^{\text{sc},\theta}$ is the complex-scaled scattered part of the wave function, $\sigma_{\ell}$ is the Coulomb phase shift, $F_{\ell}$ is the regular Coulomb function, $\eta$ is the Sommerfeld parameter, $k=\sqrt{2\mu E}$ is the wave number, and $\tilde{U}_{N}$ is the modified interaction defined as:
@@
where $z_{1}$ and $z_{2}$ are the charge numbers of the two nuclei. For a more detailed derivation of the complex scaling method for solving scattering problems, one may refer to Ref. [13], in which we have developed a computer code, COLOSS, to solve the two-body scattering problem using the complex scaling method. Since $\psi_{\ell}^{\text{sc},\theta}$ is exponentially decaying after the complex scaling operation, we are able to use a complete set of square-integrable functions as the variational basis to expand it. For a direct solution of the complex scaling method for a given parameter set of the Schrödinger equation, which can be called the exact solution, we expand $\psi_{\ell}^{\text{sc},\theta}$ with $x$-regularized Lagrange-Laguerre functions [16]:
$\psi_{\ell}^{\text{sc},\theta}(r)=\sum_{j}^{N_{r}}c_{j}f_{j}(r),$ (7)
where $c_{j}$ represents the expansion coefficient, $f_{j}(r)$ denotes the $x$-regularized Lagrange-Laguerre function, and $N_{r}$ is the number of basis functions. By substituting Eq. (7) into Eq. (5) and projecting on $f_{i}$, the inhomogeneous ordinary differential equation can be transformed into a set of linear equations:
@@
$\mathbf{H}^{\theta}_{ij}=\int\mathrm{d}r\,f_{i}(r)H^{\theta}(r;\bm{\Omega})f_{j}(r),$ (9)
@@
$b_{i}^{\theta}(\bm{\Omega})=\int\mathrm{d}r\,f_{i}(r)e^{i\theta/2}e^{i\sigma_{\ell}}\tilde{U}_{N}(re^{i\theta};\bm{\Omega})F_{\ell}(\eta,kre^{i\theta}).$ (10)
@@
$\phi_{k}(r)=\sum_{i=1}^{N_{r}}x^{(k)}_{i}f_{i}(r),$ (12)
where $\phi_{k}$ represents the $k$th reduced basis, and $x^{(k)}_{i}$ denotes the $i$th expansion coefficient of the $k$th principal component vector. All the Lagrange functions in Eq. (7), (9), and (10) can be replaced by $\{\phi_{k}\}_{k=1}^{n}$ as the new basis, and the approximated linear equation, similar to Eq. (8), to be solved for emulation can be derived. In another word, the emulator approximates the linear equation for a given target parameter set via projection onto the reduced basis, facilitating a straightforward solution. Consequently, the number of basis functions can be reduced from $N_{r}$ to $n$, leading to a reduced time complexity from $O(N_{r}^{3})$ to $O(n^{3})$ for solving the linear equation.
@@ 3.1 Application to $n$+40Ca elastic scattering
We apply our emulator first to $n$+40Ca elastic scattering at $E=20$ MeV in the center-of-mass frame. We include partial-wave channels with angular momentum $\ell\leq 10$ to ensure convergence.
@@
To construct the reduced basis for our emulator, we initially select 100 parameter sets, $\{\bm{\Omega}_{i}\}_{i=1}^{100}$, employing Latin hypercube sampling (LHS) [18] to generate values within a ±$\pm 20\%$ range centered on the optimal values from the Koning-Delaroche (KD) parameterization [19]. The angular momentum quantum numbers, $\ell$, are chosen as random integers from 0 to 10. We perform exact calculations for these parameter sets, yielding 100 exact solutions. Subsequently, principal component analysis (PCA) is applied to these solutions to derive reduced bases of 10, 12, and 14 dimensions, which are used to train our emulator.
Given the stochastic nature of the parameter generation in the training procedure, the emulator outputs are inherently random. Therefore, to evaluate the numerical accuracy and stability of our method more robustly, it is essential to perform multiple emulations. We conduct the emulation process 20 times to observe the variability and reliability of the errors. The mean absolute relative error, computed as follows, serves as a criterion for assessing our method:
$\epsilon=\sum_{i}\frac{|\sigma_{\text{E},i}-\sigma_{0}|}{\sigma_{0}},$ (13)
@@
where $\epsilon$ represents the mean of absolute relative errors, $\sigma_{\text{E},i}$ denotes the $i$th emulated cross-section, and $\sigma_{0}$ is the high-fidelity result. By averaging the errors over multiple emulations, we can obtain a more reliable measure of the emulator’s performance, thus accounting for the inherent randomness in the training parameters and providing a robust evaluation of numerical accuracy and stability.
Figure 1 shows (a) the exact and emulated differential cross sections as a function of the scattering angle and (b) their corresponding mean absolute relative errors for several sizes of the reduced basis as a function of the scattering angle.
@@
Figure 1: (a) Exact and emulated elastic scattering differential cross sections of the $n$+40Ca reaction at $E_{\text{c.m.}}=20$ MeV as a function of scattering angles. (b) Mean absolute relative errors for several sizes of the reduced basis as a function of scattering angles. See text for more details.
@@
In panel (a), the emulation results with a reduced basis of 14 dimensions are depicted by red dashed lines, while the exact results are shown using black solid lines. The emulator effectively replicates the exact angular distribution of the cross section, demonstrating its accuracy. In panel (b), the mean absolute relative errors associated with reduced bases of 10, 12, and 14 dimensions are indicated by green dotted, blue dot-dashed, and red dashed lines, respectively. The figure illustrates that the accuracy of the emulator improves approximately by an order of magnitude with each increment of two in the size of the reduced basis, highlighting the emulator’s high accuracy and rapid convergence. Notably, error peaks are observed at angles where the cross section is minimal and hard to measure accurately. This observation underscores that, despite the extensive parameter space, our emulator can deliver highly precise solutions within a considerably smaller subspace, utilizing a single set of reduced bases across various angular momentum channels.
@@ 3.2 Application to 11Be+64Zn elastic scattering
We select 200 sets of training parameters $\{\bm{\Omega}_{i}\}_{i=1}^{200}$. The potential parameters are randomly generated using LHS within a ±$\pm 20\%$ interval of the best-fit values of KD02 [19]. The angular momentum quantum numbers, $\ell$, are random integers between 0 and 60. PCA is then performed on the solutions of these parameter sets to obtain 20, 25, and 30 reduced bases for training the emulator. The emulation process is repeated 10 times, and the mean of the absolute relative errors with respect to the exact solutions, as defined in Eq. (13), is evaluated. Figure 2 (a) shows the simulated and emulated differential cross sections as a function of the scattering angle, along with the experimental data taken from Ref. [20]. Figure 2 (b) displays the corresponding mean of absolute relative errors for various sizes of the reduced basis as a function of the scattering angle.
Figure 2: (a) Exact and emulated elastic scattering cross section of 11Be+64Zn reaction at $E_{\text{c.m.}}=24.5$ MeV. The data taken from Ref.[20] is also included. (b) Mean of absolute relative error with respect to the exact results for several sizes of reduced basis. See text for more details.
As depicted in panel (a) of Figure 2, the exact results (black solid line) and the emulation results with a reduced basis of 30 dimensions (red dashed line) are indistinguishable, with both lines perfectly overlapping and accurately reproducing the experimental data (yellow circles). This outcome confirms that the complex scaling-based emulator can effectively emulate across a wide range of angular momentum values. In panel (b), the mean of absolute relative errors associated with reduced basis sizes of 20, 25, and 30 are represented by green dotted, blue dot-dashed, and red dashed lines, respectively. A clear trend of rapid convergence is evident; the error diminishes substantially as the basis size increases. Notably, at scattering angles up to approximately $70^{\circ}$ — where data are available — the relative errors remain well around $10^{-3}$ for the 20 reduced basis scenario, significantly surpassing the typical experimental uncertainty of 5%. These results underscore our emulator’s high efficiency and accuracy in simultaneously emulating across various channels and potential parameters with a single set of reduced bases. This demonstrates its substantial practicality for handling realistic scenarios in nuclear reactions research.

### v2

@@ 1 Introduction
Nuclear reactions play a crucial role in helping us understand how nuclei are formed and the properties of dense nuclear matter, like that found in neutron stars [1]. These reactions are fundamental to many astrophysical events, especially those happening in stars and during explosive occurrences like supernovae. This makes these reactions very important for astrophysics.
However, the models we use to understand these reactions depend heavily on certain assumptions, like the optical model potential, which is a way to describe how particles interact. These models need parameters that are adjusted to fit the observed data, such as elastic scattering cross sections. This fitting process introduces many uncertainties into our analysis of nuclear reactions. To deal with these uncertainties, it is vital to use robust methods for error quantification, such as Bayesian analysis, to make our results more reliable and understandable [2, 3].
@@
Several types of nuclear scattering emulators have been developed. For example, emulators based on the Kohn variational principle [6, 7], other variational methods [8, 9], and the $R$-matrix theory [10, 11] have shown impressive accuracy and efficiency. These emulators can quickly provide results that are very close to those obtained from detailed and time-consuming calculations.
However, there are some limitations. One major issue is that most emulators are designed to work with specific angular momentum states, which means they cannot easily be used for different angular momentum settings. This specialization increases the amount of storage needed and the computational time required, as the unique properties of each angular momentum state are not fully utilized. This limitation reduces the overall utility of these emulators, highlighting the need for more flexible and general approaches.
@@
The paper is organized as follows: In Sec. 2, we provide a brief overview of the complex scaling method and the development of our scattering emulator. Sec. 3 presents some applications of our emulator, and an anomaly detection test. Finally, we conclude with a summary and discuss future research directions in Sec. 4.
@@ 2 Formalism
$U_{N}(r;\bm{\omega})=$ $V_{C}(r;R_{c})+V_{0}\mathscr{Y}\left(r;R_{R},a_{R}\right)$ (2)
$+iW_{0}\mathscr{Y}\left(r;R_{W},a_{W}\right)+iW_{S}\frac{\mathrm{d}}{\mathrm{d}r}\mathscr{Y}\left(r;R_{ws},a_{ws}\right),$
where $\mathscr{Y}(r;R,a)\equiv(1+\mathrm{exp}((r-R)/a))^{-1}$ is the Woods-Saxon (WS) function, $V_{C}$ is the Coulomb interaction, the second and third terms are the real and imaginary volume terms, and the last term is the imaginary surface term. The parameter configuration is:
@@
where $\bm{\Omega}=\{\bm{\omega},\ell\}$ is the total parameter configuration we consider in this work. It should be noted that in some conventional methods for solving Eq. (1), the angular momentum quantum number $\ell$ serves not only as a parameter in the Hamiltonian but also specifies the boundary condition of the scattering wave function. This presents significant challenges when dealing with states of different angular momentum. To remove the dependence on different boundary conditions and treat $\ell$ just as a parameter, we apply the complex scaling method (CS) to solve Eq. (1). In the complex scaling method, the wave function is first separated into two parts
$\psi_{\ell}(r)=e^{i\sigma_{\ell}}F_{\ell}(\eta,kr)+\psi_{\ell}^{\text{sc}}(r),$ (5)
where $\sigma_{\ell}$ is the Coulomb phase shift, $F_{\ell}$ is the regular Coulomb function, $\eta$ is the Sommerfeld parameter, $k=\sqrt{2\mu E}/\hbar$ is the wave number, and $\psi_{\ell}^{\text{sc}}(r)$ is referred to as the scattered part of the wave function. The coordinate of $\psi_{\ell}^{\text{sc}}(r)$ is then rotated by an angle $\theta$ to get the complex-scaled scattered part of the wave function
$\psi_{\ell}^{\text{sc},\theta}(r)=e^{i\theta/2}\psi_{\ell}^{\text{sc}}(re^{i\theta}),$ (6)
and it satisfies the following inhomogeneous equation:
@@
where $H^{\theta}$ is the complex-scaled Hamiltonian, and $\tilde{U}_{N}$ is the modified interaction defined as:
@@
where $z_{1}$ and $z_{2}$ are the charge numbers of the two nuclei. Upon separating the wave function, the partial wave expanded nuclear scattering amplitude can be decomposed into two terms as well:
$f_{\ell}=f_{\ell}^{\text{Born}}+f_{\ell}^{\text{sc}},$ (9)
where the Born term of the scattering amplitude is defined as:
[…]
and the correction to the Born term of the scattering amplitude can be derived with the complex-scaled scattered part of the wave function using Cauchy theorem:
[…]
The total nuclear scattering amplitude, denoted as $\mathscr{F}_{N}$, can be obtained as a sum over the partial wave expanded amplitudes:
$\mathscr{F}_{N}(\cos{\theta})=\sum_{\ell}(2\ell+1)e^{2i\sigma_{\ell}}f_{\ell}P_{\ell}(\cos{\theta}),$ (12)
where $P_{\ell}$ is the Legendre polynomial. The angle distribution of the cross section is given by the modular square of the total scattering amplitude:
$\frac{d\sigma}{d\Omega}=\left|\mathscr{F}_{N}+\mathscr{F}_{\mathrm{C}}\right|^{2},$ (13)
where $\mathscr{F}_{C}$ is the Coulomb scattering amplitude. For a more detailed derivation of the complex scaling method for solving scattering problems and the calculation of the cross section, one may refer to Ref. [13], in which we have developed a computer code, COLOSS, to solve the two-body scattering problem using the complex scaling method.
Since $\psi_{\ell}^{\text{sc},\theta}$ is exponentially decaying after the complex scaling operation, we are able to use a complete set of square-integrable functions as the variational basis to expand it. For a direct solution of the complex scaling method for a given parameter set of the Schrödinger equation, which can be called the exact solution, we expand $\psi_{\ell}^{\text{sc},\theta}$ with $x$-regularized Lagrange-Laguerre functions [16]:
$\psi_{\ell}^{\text{sc},\theta}(r)=\sum_{j}^{N_{r}}c_{j}g_{j}(r),$ (14)
where $c_{j}$ represents the expansion coefficient, $g_{j}(r)$ denotes the $x$-regularized Lagrange-Laguerre function, and $N_{r}$ is the number of basis functions. By substituting Eq. (14) into Eq. (7) and projecting on $g_{i}$, the inhomogeneous ordinary differential equation can be transformed into a set of linear equations:
@@
$\mathbf{H}^{\theta}_{ij}=\int\mathrm{d}r\,g_{i}(r)H^{\theta}(r;\bm{\Omega})g_{j}(r),$ (16)
@@
$b_{i}^{\theta}(\bm{\Omega})=\int\mathrm{d}r\,g_{i}(r)e^{i\theta/2}e^{i\sigma_{\ell}}\tilde{U}_{N}(re^{i\theta};\bm{\Omega})F_{\ell}(\eta,kre^{i\theta}).$ (17)
With the solution of the linear equation, the correction to the Born term of the scattering amplitude can be simplified as [13]:
$f^{\text{sc}}_{\ell}=-\frac{e^{-2i\sigma_{\ell}}}{E}\sum_{i}c_{i}(\bm{\Omega})b_{i}^{\theta}(\bm{\Omega}).$ (18)
@@
$\phi_{k}(r)=\sum_{i=1}^{N_{r}}x^{(k)}_{i}g_{i}(r),$ (20)
where $\phi_{k}$ represents the $k$th reduced basis, and $x^{(k)}_{i}$ denotes the $i$th expansion coefficient of the $k$th principal component vector. All the Lagrange functions in Eq. (14), (16), and (17) can then be replaced by $\{\phi_{k}\}_{k=1}^{n}$ as the new basis, and the approximated linear equation, similar to Eq. (15), to be solved for emulation can be derived. The scattering amplitude can be emulated according to Eq. (18) with the solutions of the approximated linear equation and its inhomogeneous terms. In other words, the emulator approximates the linear equation for a given target parameter set via projection onto the reduced basis, facilitating a straightforward solution. Consequently, the number of basis functions can be reduced from $N_{r}$ to $n$, leading to a reduced time complexity from $O(N_{r}^{3})$ to $O(n^{3})$ for solving the linear equation.
@@ 3.1 Application to $n$+40Ca elastic scattering
We apply our emulator first to $n+^{40}\text{Ca}$ elastic scattering at $E=20$ MeV in the center-of-mass frame. We include partial-wave channels with angular momentum $\ell\leq 10$ to ensure convergence.
@@
To construct the reduced bases for our emulator, we initially select 100 parameter sets, $\{\bm{\Omega}_{i}\}_{i=1}^{100}$, randomly generated within a ±$\pm 20\%$ range centered on the optimal values derived from the Koning-Delaroche (KD) parameterization [18]. We perform this selection using Latin hypercube sampling (LHS) [19], and we confine our parameter space to this range. Angular momentum quantum numbers, $\ell$, are randomly chosen as integers from 0 to 10.111Including parameter sets with low angular momentum in the training set is essential, while excluding higher $\ell$ values does not significantly reduce accuracy. This is because the centrifugal barrier term for states with high angular momentum dominates over the optical potential, making it challenging for the emulator to capture the analytic properties of the optical potential effectively. We perform exact calculations for these training parameter sets, yielding 100 exact solutions. Subsequently, principal component analysis (PCA) is applied to these solutions to derive reduced bases of 10, 12, and 14 dimensions, which are used to train our emulator.
With these reduced bases, we first emulate the cross section on the parameters provided by the Koning-Delaroche (KD) parameterization [18] to make a comparison with the exact results. To robustly assess the numerical accuracy and stability of our method, it is crucial to conduct multiple emulations. Consequently, we generated another 100 sets of target parameters across the entire parameter space using the Latin Hypercube Sampling (LHS) method and repeated the emulation on these sets to examine the variability and reliability of the errors. The mean absolute relative error of 100 results, computed as follows, serves as a criterion for assessing our method:
$\epsilon=\sum_{i}\frac{|\sigma_{\text{E},i}-\sigma_{0,i}|}{\sigma_{0,i}},$ (21)
@@
where $\epsilon$ represents the mean of absolute relative errors, and $\sigma_{\text{E},i}$ and $\sigma_{0,i}$ denote the emulated and high-fidelity cross-section for the $i$th target parameter set. By averaging the errors over multiple emulations, we can obtain a more reliable measure of the emulator’s performance and provide a robust evaluation of numerical accuracy and stability. The results are shown in Fig. 1.
@@
Figure 1: (a) Exact and emulated elastic scattering differential cross sections of the $n+^{40}\text{Ca}$ reaction at $E_{\text{c.m.}}=20$ MeV as a function of scattering angles. (b) Mean absolute relative errors of 100 sets of target parameters for several sizes of the reduced bases as a function of scattering angles. See text for more details.
In panel (a) of Fig. 1, the exact and emulated differential cross sections as a function of the scattering angle are illustrated by red dashed and black solid lines. Values of the parameters are taken from the Koning-Delaroche parameterization  [18]. We have adopted 14 reduced basis states for the emulation and 60 $x$-regularized Lagrange-Laguerre basis for the exact calculation. It can be observed that our emulator effectively replicates the angular distribution of the cross section, demonstrating the emulator’s accuracy.
@@
In panel (b) of Fig. 1, the mean absolute relative errors for 100 sets of target parameters are plotted as a function of the scattering angle for several reduced basis sizes. The errors associated with reduced bases of 10, 12, and 14 dimensions are represented by green dotted, blue dot-dashed, and red dashed lines, respectively. The figure demonstrates that the accuracy of the emulator significantly improves—approximately by an order of magnitude—with each two-dimensional increment in the reduced basis size, underscoring the emulator’s rapid convergence and high accuracy. Notably, error peaks occur at angles where the cross section is minimal and challenging to measure accurately. This highlights the emulator’s capability to deliver highly precise solutions across a broad parameter space, even while operating within a considerably smaller subspace and using a consistent set of reduced bases across different angular momentum channels. With the performance testing across the entire parameter space, we can estimate the uncertainty introduced by the emulator by assuming it has a constant relative error, as detailed in Ref. [3].
@@ 3.2 Application to 11Be+64Zn elastic scattering
To construct the reduced basis for the emulator, we initially selected 200 sets of training parameters, denoted as $\{\bm{\Omega}_{i}\}_{i=1}^{200}$. These potential parameters were randomly generated using Latin Hypercube Sampling (LHS) within a ±$\pm 20\%$ interval of the best-fit values reported in Ref. [20], with the parameter space confined to this range. The angular momentum quantum numbers, $\ell$, were chosen as random integers between 0 and 60. Principal Component Analysis (PCA) was subsequently applied to the solutions of these parameter sets to derive 20, 25, and 30 reduced bases for training the emulator.
We initially applied the emulator to the best-fit parameters from Ref. [20] to facilitate a direct comparison with both the exact results and the experimental data. Following this initial test, the emulator’s performance was evaluated on 100 sets of target parameters, also randomly generated using LHS across the entire parameter space. The mean absolute relative errors with respect to the exact solutions, as specified in Eq. (21), were then calculated. The results of these tests are presented in Fig. 2.
Figure 2: (a) Exact and emulated elastic scattering cross section of 11Be+64Zn reaction at $E_{\text{c.m.}}=24.5$ MeV. The data taken from Ref. [20] is also included. (b) Mean absolute relative errors of 100 sets of target parameters with respect to the exact results for several sizes of reduced basis. See text for more details.
Figure 2 (a) displays the differential cross sections as a function of the scattering angle, comparing exact, emulated, and experimental results. For this analysis, 30 reduced basis states were utilized for the emulation, while 100 $x$-regularized Lagrange-Laguerre basis were employed for the exact calculations. In this panel, the exact results (black solid line) and the emulation results (red dashed line) are indistinguishable, with both lines perfectly overlapping and accurately reproducing the experimental data (yellow circles). This outcome confirms that the complex scaling-based emulator can effectively emulate across a wide range of angular momentum values.
Figure 2 (b) displays the mean of absolute relative errors of 100 different sets of parameters for various sizes of the reduced basis. The errors are plotted as a function of the scattering angle, with different basis sizes represented by distinct line styles: 20 basis states by a green dotted line, 25 basis states by a blue dot-dashed line, and 30 basis states by a red dashed line. A clear trend of rapid convergence is evident; the error diminishes substantially as the basis size increases. Notably, at scattering angles up to approximately $70^{\circ}$ — where data are available — the relative errors remain well around $10^{-3}$ for the 20 reduced basis scenario, significantly surpassing the typical experimental uncertainty of 5%. These results underscore our emulator’s high efficiency and accuracy in simultaneously emulating across various channels and potential parameters across the entire parameter space with a single set of reduced bases. This demonstrates its substantial practicality for handling realistic scenarios in nuclear reactions research.